Suppose the sum of two numbers is 16. Find the two such numbers such that the sum of their squares would be a minimum.
Answer by Cintchr(475)

1 + 15 = 16 ..... 1^2 + 15^2 = 1 + 225 = 226
2 + 14 = 16 ..... 2^2 + 14^2 = 4 + 196 = 200
and so on and so on ....
8 + 8 = 16 ...... 8^2 + 8^2 = 64 + 64 = 128
This is your minimum